A merchant blends tea that sells for $4.10 a pound with tea that sells for $2.10 a pound to produce 60 lb of a mixture that sells for $3.30 a pound. How many pounds of each type of tea does the merchant use in the blend?
I have tried setting this problem up so many different ways and continue to get it wrong. Do you know how it should be set up to solve? Any help would be appreciated. Thank you!!

Answer by ewatrrr(24785)

 
Hi
Traditionally I will assign variable x to the more expensive (or higher concentration, in the case of %)
4.10x + 2.10(60-x) = 3.30(60)
x = 1.20(60)/2 |Note: doing it this way makes the Arithmetic easier 4.10-2.10+=+2